Banks Management

The following statistics and estimates were compiled by Big Moon Bank regarding a proposed new branch office and the bank itself:

Branch office expected return                                                                        = 15%

            Standard deviation of branch return                                                   = 8%

            Existing bank’s expected return                                                         = 10%

            Standard deviation of existing bank’s return                                     = 5%

            Branch asset value as a percentage of total bank assets                     = 16%

            Correlation of net cash flows for branch and bank as a whole           = +0.48

 

What will happen to Big Moon’s total expected return and overall risk if the proposed new branch project is adopted?  
Answer #1

if the project is accepted, the total expected return and overall risk becomes 10.8 % and 4.94 %, respectively.


TOTAL EXPECTED RETURN =0.16 × 15+ 0.84 × 10 =2.4+8.4 =10.8 %


 OVERALL RISK =0.16² × 8²+0.84² × 5²+ 2 × 0.16 × 0.84 × 8 × 5 × 0.48 =0.0256 × 64+0.7056 × 25+2 × 0.1344 x 40 × 0.48 =1.6384+17.64+5.16096 =4.94 %


answered by: deanmoreno